# Hessian Aware Low-Rank Weight Perturbation for Continual Learning

### v2

@@ 4.2 Convolutional Layers
In addition, we can also decompose the weights of a convolutional layer in a similar way. Suppose that the size of the convolutional kernel is $d\times d$. The base weights of the convolution layer for task $\mathcal{T}_{0}$ would be a tensor $\mathbf{W}_{\text{conv}}^{\text{base}}\in\mathbb{R}^{d\times d\times J\times I}$. Similar to Eq. 2, a low-rank weight perturbation for transforming $\mathbf{W}_{\text{conv}}^{\text{base}}$ to $\mathbf{W}_{\text{conv},1}\in\mathbb{R}^{d\times d\times J\times I}$ is,
@@
where $\mathbf{R}_{\text{conv},1}\in\mathbb{R}^{1\times 1\times J\times 1}$, $\mathbf{S}_{\text{conv},1}\in\mathbb{R}^{1\times 1\times 1\times I}$ and $\mathbf{B}_{\text{conv},1}\in\mathbb{R}^{1\times 1\times J\times I}$ is sparse low-rank tensor (matrix), $\otimes$ and $\oplus$ are element-wise tensor multiplication and summation operators will automatically expand tensors to be of equal sizes, following the broadcasting semantics of some popular scientific computation package like Numpy [36] or PyTorch [37]. Thus, the number of parameters added through this kind of decomposition is still $\mathcal{O}(I+J)$.
To get the estimations of the introduced parameters, we can solve a similar LSE problem as in Eq. 3 to obtain initial estimates $\mathbf{R}_{\text{conv},1}^{\text{free}}$ and $\mathbf{S}_{\text{conv},1}^{\text{free}}$. Then we take the average of the first two dimensions to transform the discrepancy tensor $(\mathbf{W}_{\text{conv},1}^{\text{free}}-\mathbf{R}_{\text{conv},1}^{\text{free}}\otimes\mathbf{W}_{\text{conv}}^{\text{base}}\otimes\mathbf{S}_{\text{conv},1}^{\text{free}})$ to be a $1\times 1\times J\times I$ tensor, which can then be applied a similar decomposition with the FC layers (Eq. 5 and Eq. 6.) to obtain the low-rank estimates of $\mathbf{B}_{\text{conv},1}^{\text{free}}$.
@@ 4.3 Rank Selection for each layer based on Hessian Information
Assume that a neural network of $L$ layers with vectorized weights $(\boldsymbol{\omega}^{\star}_{1},\dots,\boldsymbol{\omega}^{\star}_{L})$ that have converged to local optima, such that the first and second order optimally conditions are satisfied, i.e., the gradient is zero, and the Hessian is positive semi-definite. Suppose a perturbation $\Delta\boldsymbol{\omega}^{\star}_{1}$ applied to the first layer weights, then we have the loss increment
$\begin{split}\mathcal{L}(\boldsymbol{\omega}^{\star}_{1}-\Delta\boldsymbol{\omega}^{\star}_{1},&\dots,\boldsymbol{\omega}^{\star}_{L})-\mathcal{L}(\boldsymbol{\omega}^{\star}_{1},\dots,\boldsymbol{\omega}^{\star}_{L})\\ &\leq\|\frac{1}{2}\mathbf{H}_{1}\|_{F}\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F}+o(\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F}),\end{split}$ (8)
where $\mathbf{H}_{1}=\Delta\mathcal{L}(\boldsymbol{\omega}^{\star}_{1})$ is the Hessian matrix at only the variables of the first layer weights.
@@
In our low-rank perturbation setting, we assume that $\mathbf{W}_{1}^{\text{free}}$ by warm-up training is a local optimum. By Theorem 1, we consider the difference between $\mathbf{W}_{1}^{\text{free}}$ and its $k$-rank approximation $\mathbf{W}_{1}^{(k){\text{free}}}$ as a perturbation $\Delta\mathbf{W}_{1}^{(k){\text{free}}}$ for the weights. Then the amount of perturbation can be computed with the low-rank approximation error:
$\begin{split}\|\Delta\mathbf{W}_{1}^{(k){\text{free}}}\|_{F}=&\|\mathbf{W}_{1}^{\text{free}}-\mathbf{W}_{1}^{(k){\text{free}}}\|_{F}=\sqrt{\sum_{i=k+1}^{r}\sigma_{i}^{2}}\\ \end{split}$ (9)
@@
$\begin{split}\mathcal{L}(\mathbf{W}_{1}^{\text{free}}&-\Delta\mathbf{W}_{1}^{(k){\text{free}}})-\mathcal{L}(\mathbf{W}_{1}^{\text{free}})\\ &\leq\|\frac{1}{2}\mathbf{H}_{1}\|_{F}(\sum_{i=1}^{k}\sigma_{i}^{2})+o(\sum_{i=k+1}^{r}\sigma_{i}^{2}).\end{split}$ (10)
where the Hessian matrix $\mathbf{H}_{1}$ can be approximated by the negative empirical Fisher information [40], i.e., the outer product of the gradient vector for the layer weights. So $\|\mathbf{H}_{1}\|_{F}$ can be approximated by $\|\mathbf{g}_{1}\|_{2}^{2}$, where $\mathbf{g}_{1}=\frac{\partial\mathcal{L}}{\partial W_{1}}\mid_{W_{1}=W_{1}^{\text{free}}}$. Finally, we can quantitatively measure the contribution of the loss of adding a marginal rank $k$ for a particular layer $l$ by $\|\mathbf{g}_{l}\|_{2}^{2}\sigma_{l,k}^{2}$, where $\mathbf{g}_{l}$ is the gradient for the layer-$l$ weights and $\sigma_{l,k}$ is the $k$-th singular value of the free-trained layer-$l$ weights, and sort them by the descending order of importance.
For a given loss approximation rate $\alpha$ (e.g., 0.9), we can determine the rank $k_{l}$ (with $k_{l}\leq r_{l}$ where $r_{l}$ is the total rank of the layer $l$) for each layer $l=0,...,L-1$ by solving
$\min_{k_{1},...,k_{L-1}}\quad\sum_{l=0}^{L-1}\sum_{i=1}^{k_{l}}\|\mathbf{g}_{l}\|_{2}^{2}\sigma_{l,i}^{2}$ (11)
@@
4.4 Regularization and Pruning on the Introduced Parameters
@@
$\mathcal{L}_{\text{reg}}(\mathbf{W}_{t})=\sum_{l}$ $\Big{[}\lambda_{0}(\|\mathbf{U}_{t,l}^{(k_{l})\text{free}}\|+\|\mathbf{V}_{t,l}^{(k_{l})\text{free}}\|)$ (12)
@@
where $\lambda_{0},\lambda_{1}$ are balacing coefficients, and the subscripts $(t,l)$ indicate the revelant weights for layer $l$ in task $t$.
@@
0:  Task data $\{\mathcal{D}_{t}\}_{i=1}^{T}$; total epochs for one task $n$; rank estimation epochs $n_{r}$; parameter increments limitation ratio $p$, approximation rate $\alpha$.
@@
3:     Warm-up pre-training on task $\mathcal{T}_{t}$ for $n_{r}$ epochs: $\mathbf{W}^{\text{free}}_{t}=\arg\min_{\mathbf{W}}L(\mathbf{W};\mathcal{D}_{t})$.
@@
5:     Apply $\mathbf{U}_{t}^{(k)\text{free}},\boldsymbol{\Sigma}_{t}^{(k)\text{free}},\mathbf{V}_{t}^{(k)\text{free}}=\texttt{SVD}(\mathbf{B}^{\text{free}}_{t})$.
@@
$\mathbf{W}_{t}^{\star}=\arg\min_{\mathbf{W}}\mathcal{L}(\mathbf{W};\mathcal{D}_{t})+\mathcal{L}_{\text{reg}}(\mathbf{W})$ (13)
@@
4.5 Summary of Algorithm
In the previous sections, we described the model update from $\mathcal{T}_{0}$ to $\mathcal{T}_{1}$ with an illustrative example. The overall description of our proposed HALRP is presented in Algorithm 1.
At the start, the learner was trivially trained on the initial task $\mathcal{T}_{0}$ to obtain the base weights $\mathbf{W}^{\text{base}}$. As for each incoming task $\mathcal{T}_{t}$ with $t=1,...,T-1$, we first train the model without any constraints for $n_{r}$ epochs to get a rough initialization $\mathbf{W}^{\text{free}}_{t}$. Secondly, we apply the low-rank decomposition on all layers with Eq. 2 (for linear layers) for Eq. 7 (for convolutional layers) by solving a least square error minimization problem described in Eq. 4. Then, we further apply the singular value decomposition for the residual matrix $\mathbf{B}^{\text{free}}_{t}$. With this decomposition, we can measure the Hessian-aware perturbations and the rank $k_{l}$ for each layer $l$ through Eq. 11. After the rank selection, we can re-initialize the model parameters with the approximated weights $\mathbf{W}^{(k)\text{free}}_{t}\approx\mathbf{W}^{\text{free}}_{t}$ and then fine-tune the model for the remaining $n-n_{r}$ epochs to obtain the optimal weights for $\mathcal{T}_{t}$. As for the inference stage, the base weights and the task-specific parameters can be adopted to make predictions for each task.
@@ 5 Experimental Results
CIFAR-100 Split (with LeNet) CIFAR-100 SuperClass (with LeNet)
@@
5.1 Experimental Settings
Datasets: We evaluate the algorithm on the following datasets: 1) CIFAR-100 Split: Split the classes into 10 groups; for each group, consider a 10-way classification task. 2) CIFAR-100-Superclass: It consists of images from 20 superclasses of the CIFAR-100 dataset. 3) Permuted MNIST (P-MNIST): obtained from the MNIST dataset by random permutations of the original MNIST pixels. We follow [41] to create 10 sequential tasks using different permutations, and each task has 10 classes. 4) Five-dataset: It uses a sequence of 5 different benchmarks including CIFAR10 [42], MNIST [43], notMNIST [44], FashionMNIST [45] and SVHN [46]. Each benchmark contains 10 classes. 5) Omniglot Rotation [47]: 100 12-way classification tasks. The rotated images in $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$ are generated by following [9]. 6) TinyImageNet: a variant of ImageNet [48] dataset containing 200 classes. Here, we adopted two settings, one with 20 10-way classification tasks (TiyImageNet 20-split) and another more challenging setting with 40 5-way classification tasks (TinyImageNet 40-split).
Baselines: We compared the following baselines by the publicly released code or our re-implementation: STL: single task learning with individual models for each task. MTL: multi-task learning with a single for all the tasks simultaneously [49]. EWCElastic Weight Consolidation method proposed by [4]L2: the model is trained with $L_{2}$-regularizer [4] $\lambda\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t-1}\|_{2}^{2}$ between the current model and the previous one. BN: The Batch Normalization method [50]. BE: The Batch Ensemble method proposed by [11]. APD: The Additive Parameter Decomposition method [9]. Each layer of the target network was decomposed into task-shared and task-specific parameters with mask vectors. APDfix: We modify the APD method by fixing the model parameters while only learning the mask vector when a new task comes to the learner. IBPWF: Determine the model expansion with non-parametric Bayes and weights factorization [10]. GPM [2]: A replay-based method by orthogonal gradient descent. WSN [19]: Introduce learnable weight scores to generate task-specific binary masks for optimal subnetwork selection. BMKP [51]: A bilevel memory framework for knowledge projection: a working memory to ensure plasticity and a long-term memory to guarantee stability. Additionally, we also compared our methods with IRU [33], a method also based on the low-rank decomposition. However, due to the code limitation [^5] IRU code only contains the implementation on the multi-layer perceptron. See https://github.com/CSIPlab/task-increment-rank-update of IRU [33], we further implemented our method under the dataset protocol and model architecture setting in [33] and compared our performance with the results reported in [33]. For the implementation of all the methods, we applied same hyperparameters (e.g., batch size, training epochs, regularization coefficient) to realize fair comparisons. See Appendix F.3 for more details.
Model Architecture: For CIFAR-100 Split, CIFAR-100 Superclass, and P-MNIST datasets, we adopted LeNet as the base model. As for Five-dataset, we tested with AlexNet and reduced ResNet18 networks wehre the latter has reduced filters standard Resnet18. And we followed [19] to use an extended LeNet model on the Omniglot-Rotation dataset. As for TinyImageNet dataset, we also adopted two architectures: AlexNet and reduced ResNet18. See Appendix F.2 for more details.
Evaluation Metrics: We mainly adopted the average of the accuracies of the final model on all tasks (we simply call “accuracy” or “Acc.” later) for the empirical comparisons. As for the forgetting statistics, we applied backward transfer (BWT) [2] as a quantitive measure. For each single experiment (e.g., each task order), we repeat with five random seeds to compute the average and standard error.

### v5

@@ 4.2 Convolutional Layers
In addition, we can decompose the weights of a convolutional layer in a similar way. Suppose that the size of the convolutional kernel is $d\times d$. The base weights of the convolution layer for task $\mathcal{T}_{0}$ would be a tensor $\mathbf{W}_{\text{conv}}^{\text{base}}\in\mathbb{R}^{d\times d\times J\times I}$. Similar to Eq. 2, a low-rank weight perturbation for transforming $\mathbf{W}_{\text{conv}}^{\text{base}}$ to $\mathbf{W}_{\text{conv},1}\in\mathbb{R}^{d\times d\times J\times I}$ is,
@@
where $\mathbf{R}_{\text{conv},1}\in\mathbb{R}^{1\times 1\times J\times 1}$, $\mathbf{S}_{\text{conv},1}\in\mathbb{R}^{1\times 1\times 1\times I}$ and $\mathbf{B}_{\text{conv},1}\in\mathbb{R}^{1\times 1\times J\times I}$ is sparse low-rank tensor (matrix), $\otimes$ and $\oplus$ are element-wise tensor multiplication and summation operators that will automatically expand tensors to be of equal sizes, following the broadcasting semantics of some popular scientific computation package like Numpy [36] or PyTorch [37]. Thus, the number of parameters added through this kind of decomposition is still $\mathcal{O}(I+J)$.
To get the estimations of the introduced parameters, we can solve a similar LSE problem as in Eq. 3 to obtain initial estimates $\mathbf{R}_{\text{conv},1}^{\text{free}}$ and $\mathbf{S}_{\text{conv},1}^{\text{free}}$. Then we take the average of the first two dimensions to transform the discrepancy tensor $(\mathbf{W}_{\text{conv},1}^{\text{free}}-\mathbf{R}_{\text{conv},1}^{\text{free}}\otimes\mathbf{W}_{\text{conv}}^{\text{base}}\otimes\mathbf{S}_{\text{conv},1}^{\text{free}})$ to be a $1\times 1\times J\times I$ tensor, which can then be applied a similar decomposition with the linear layers (Eq. 5 and 6.) to obtain the low-rank estimates of $\mathbf{B}_{\text{conv},1}^{\text{free}}$.
@@ 4.3 Rank Selection for each layer based on Hessian Information
Assume that a neural network of $L$ layers with vectorized weights $(\boldsymbol{\omega}^{\star}_{1},\dots,\boldsymbol{\omega}^{\star}_{L})$ that have converged to local optima, such that the first and second order optimally conditions are satisfied, i.e., the gradient is zero, and the Hessian is positive semi-definite. Suppose a perturbation $\Delta\boldsymbol{\omega}^{\star}_{1}$ applied to the first layer weights, then we have the loss change
$\begin{split}|\mathcal{L}(\boldsymbol{\omega}^{\star}_{1}-\Delta\boldsymbol{\omega}^{\star}_{1},&\dots,\boldsymbol{\omega}^{\star}_{L})-\mathcal{L}(\boldsymbol{\omega}^{\star}_{1},\dots,\boldsymbol{\omega}^{\star}_{L})|\\ &\leq\frac{1}{2}\|\mathbf{H}_{1}\|_{F}\cdot\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F}+o\left(\|\Delta\boldsymbol{\omega}^{\star}_{1}\|^{2}_{F}\right),\end{split}$ (8)
where $\mathbf{H}_{1}=\nabla^{2}\mathcal{L}(\boldsymbol{\omega}^{\star}_{1})$ is the Hessian matrix at only the variables of the first layer weights.
@@
In our low-rank perturbation setting, we assume that $\mathbf{W}_{1}^{\text{free}}$ by warm-up training is a local optimum. By Theorem 1, we consider the difference between $\mathbf{W}_{1}^{\text{free}}$ and its $k$-rank approximation $\mathbf{W}_{1}^{(k){\text{free}}}$ as a perturbation $\Delta\mathbf{W}_{1}^{\text{free}}$ for the weights. Then the amount of perturbation can be computed with the low-rank approximation error:
$\begin{split}\|\Delta\mathbf{W}_{1}^{\text{free}}\|_{F}=&\|\mathbf{W}_{1}^{\text{free}}-\mathbf{W}_{1}^{(k){\text{free}}}\|_{F}=\sqrt{\sum_{i=k+1}^{r}\sigma_{i}^{2}}\\ \end{split}$ (9)
@@
$\begin{split}|\mathcal{L}(\mathbf{W}_{1}^{(k)\text{free}})&-\mathcal{L}(\mathbf{W}_{1}^{\text{free}})|\\ &\leq\frac{1}{2}\|\mathbf{H}_{1}\|_{F}\cdot\left(\sum_{i=k+1}^{r}\sigma_{i}^{2}\right)+o\left(\sum_{i=k+1}^{r}\sigma_{i}^{2}\right)\end{split}$ (10)
where the Hessian matrix $\mathbf{H}_{1}$ can be approximated by the negative empirical Fisher information [40], i.e., the outer product of the gradient vector for the layer weights. So $\|\mathbf{H}_{1}\|_{F}$ can be approximated by $\|\mathbf{g}_{1}\|_{2}^{2}$, where $\mathbf{g}_{1}=\frac{\partial\mathcal{L}}{\partial\mathbf{W}_{1}}|_{\mathbf{W}_{1}=\mathbf{W}_{1}^{\text{free}}}$. Finally, we can quantitatively measure the contribution of the loss of adding a marginal rank $k$ for a particular layer $l$ by $\|\mathbf{g}_{l}\|_{2}^{2}\sigma_{l,k}^{2}$, where $\mathbf{g}_{l}$ is the gradient for the layer-$l$ weights and $\sigma_{l,k}$ is the $k$-th singular value of the free-trained layer-$l$ weights, and sort them by the descending order of importance.
For a given loss approximation rate $\alpha$ (e.g., 0.9), we can determine the rank $k_{l}$ (with $k_{l}\leq r_{l}$ where $r_{l}$ is the total rank of the layer $l$) for each layer $l=1,...,L$ by solving
$\min_{k_{1},...,k_{L}}\quad\sum_{l=1}^{L}\sum_{i=1}^{k_{l}}\|\mathbf{g}_{l}\|_{2}^{2}\sigma_{l,i}^{2}$ (11)
@@
4.4 Regularization and Pruning on Parameters
@@
$\mathcal{L}_{\text{reg}}(\mathbf{W}_{t};\lambda_{0},\lambda_{1})=\sum_{l}$ $\Big{[}\lambda_{0}(\|\mathbf{U}_{t,l}^{(k_{l})\text{free}}\|+\|\mathbf{V}_{t,l}^{(k_{l})\text{free}}\|)$ (12)
@@
where $\lambda_{0},\lambda_{1}$ are balancing coefficients, and the subscripts $(t,l)$ indicate the relevant weights for layer $l$ in task $t$.
@@ 4.5 Summary of Algorithm
0:  Task data $\{\mathcal{D}_{t}\}_{t=0}^{T-1}$; total epochs for one task $n$; rank estimation epochs $n_{r}$; parameter increments limitation ratio $p$, approximation rate $\alpha$.
@@
3:     Warm-up pre-training on task $\mathcal{T}_{t}$ for $n_{r}$ epochs: $\mathbf{W}^{\text{free}}_{t}=\arg\min_{\mathbf{W}}\mathcal{L}(\mathbf{W};\mathcal{D}_{t})$.
@@
5:     Apply $\mathbf{U}_{t}^{\text{free}},\boldsymbol{\Sigma}_{t}^{\text{free}},\mathbf{V}_{t}^{\text{free}}\leftarrow\texttt{SVD}(\mathbf{B}^{\text{free}}_{t})$.
@@
$\mathbf{W}_{t}^{\star}=\arg\min_{\mathbf{W}}\left[\mathcal{L}(\mathbf{W};\mathcal{D}_{t})+\mathcal{L}_{\text{reg}}(\mathbf{W};\lambda_{0},\lambda_{1})\right]$ (13)
@@
In the previous sections, we described the model update from $\mathcal{T}_{0}$ to $\mathcal{T}_{1}$ with an illustrative example. The overall description of our proposed HALRP is shown in Algorithm 1.
At the start, the learner was trivially trained on the first task $\mathcal{T}_{0}$ to obtain $\mathbf{W}^{\text{base}}$. As for each incoming task $\mathcal{T}_{t}$ with $t=1,...,T-1$, we first train the model without any constraints for $n_{r}$ epochs to get a rough initialization $\mathbf{W}^{\text{free}}_{t}$. Secondly, we apply the low-rank decomposition on all layers with Eq. 2 (for linear layers) or Eq. 7 (for convolutional layers) by solving a least square error minimization problem described in Eq. 4. Then, we further apply the singular value decomposition for the residual matrix $\mathbf{B}^{\text{free}}_{t}$. With this decomposition, we can measure the Hessian-aware perturbations and the rank $k_{l}$ for each layer $l$ through Eq. 11. After the rank selection, we can re-initialize the model parameters with the approximated weights $\mathbf{W}^{(k)\text{free}}_{t}\approx\mathbf{W}^{\text{free}}_{t}$ and then fine-tune the model for the remaining $(n-n_{r})$ epochs to obtain the optimal weights for $\mathcal{T}_{t}$. As for the inference stage, the base weights and the task-specific parameters can be adopted to make predictions for each task.
@@ 5 Experimental Results
5.1 Experimental Settings
Datasets: We evaluate the algorithm on the following datasets: 1) CIFAR100-Split: Split the classes into 10 groups; for each group, consider a 10-way classification task. 2) CIFAR100-SuperClass: It consists of images from 20 superclasses of the CIFAR-100 dataset. 3) Permuted MNIST (P-MNIST): obtained from the MNIST dataset by random permutations of the original MNIST pixels. We follow [41] to create 10 sequential tasks using different permutations, and each task has 10 classes. 4) Five-dataset: It uses a sequence of 5 different benchmarks including CIFAR10 [42], MNIST [43], notMNIST [44], FashionMNIST [45] and SVHN [46]. Each benchmark contains 10 classes. 5) Omniglot Rotation [47]: 100 12-way classification tasks. The rotated images in $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$ are generated by following [9]. 6) TinyImageNet: a variant of ImageNet [48] dataset containing 200 classes. Here, we adopted two settings, one with 20 10-way classification tasks (TiyImageNet 20-split) and another more challenging setting with 40 5-way classification tasks (TinyImageNet 40-split).
Baselines: We compared the following baselines by the publicly released code or our re-implementation: STL: single-task learning with individual models for each task. MTL: multi-task learning with a single for all the tasks simultaneously [49]. EWCElastic Weight Consolidation method proposed by [4]L2: the model is trained with $L_{2}$-regularizer [4] $\lambda\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}_{t-1}\|_{2}^{2}$ between the current model and the previous one. BN: The Batch Normalization method [50]. BE: The Batch Ensemble method proposed by [11]. APD: The Additive Parameter Decomposition method [9]. Each layer of the target network was decomposed into task-shared and task-specific parameters with mask vectors. APDfix: We modify the APD method by fixing the model parameters while only learning the mask vector when a new task comes to the learner. IBPWF: Determine the model expansion with non-parametric Bayes and weights factorization [10]. GPM [2]: A replay-based method by orthogonal gradient descent. WSN [19]: Introduce learnable weight scores to generate task-specific binary masks for optimal subnetwork selection. BMKP [51]: A bilevel memory framework for knowledge projection: a working memory to ensure plasticity and a long-term memory to guarantee stability. CLR [18]: An expansion-based method by applying Channel-Wise Linear Reprogramming transformations on each convolutional layer in the base model. This method originally relies on a model pre-trained on an extra dataset (i.e., ImageNet-1K) disjoint with the above task datasets, rather than the train-from-scratch manner adopted by other baselines. To make a fair comparison, we pre-trained the base model on the first task of the above task datasets and applied it to all tasks. PRD [52]: Prototype-sample relation distillation with supervised contrastive learning. Furthermore, we also compared our methods with IRU [34], a method also based on the low-rank decomposition. However, due to the code limitation [^5] IRU code only contains the implementation on the multi-layer perceptron. See https://github.com/CSIPlab/task-increment-rank-update of IRU [34], we further implemented our method under the dataset protocol and model architecture setting in [34] and compared our performance with the results reported in [34]. For the implementation of all the methods, we applied the same hyperparameters (e.g., batch size, training epochs, regularization coefficient) to realize fair comparisons. See Appendix F.3 for more details.
Model Architecture: For CIFAR100-Split, CIFAR100-SuperClass, and P-MNIST datasets, we adopted LeNet as the base model. As for Five-dataset and ImageNet datasets, we evaluated with AlexNet and reduced ResNet18 networks wehre the latter has reduced filters compared to standard ResNet18 (see Appendix F.2). And we followed [19] to use an extended LeNet model on the Omniglot-Rotation dataset.
Evaluation Metrics: We mainly adopted the average of the accuracies of the final model on all tasks (we simply call “accuracy” or “Acc.” later) for the empirical comparisons. As for the forgetting statistics, we applied backward transfer (BWT) [2] as a quantitive measure. Especially, we also compared the order-robustness of different methods by calculating the Order-normalized Performance Disparity that will be introduced later. For each single experiment (e.g., each task order), we repeat with five random seeds to compute the average and standard error.
CIFAR100-Split (with LeNet) CIFAR100-SuperClass (with LeNet)